UNEXPECTED GENERATION AND OBSERVATION OF A T-SHAPED COMPLEX OF H_{2}C_{2}⋯AgCCH

\underline{N. R. WALKER}, School of Chemistry, Bedson Building, Newcastle University, Newcastle-upon-Tyne, NE1 7RU, U.K.; S. L. STEPHENS, W. MIZUKAMI, D. P. TEW AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

An experiment to probe species generated within a supersonically-expanding jet consisting of SF_{6}, Ag, C_{2}H_{2} and argon by broadband rotational spectroscopy revealed the existence of a T-shaped complex of hitherto unknown origin. Empirical tests revealed that this complex requires the presence of C_{2}H_{2} and Ag within the gas sample. While the intensity of the associated transitions are enhanced by the presence of SF_{6}, theoretical calculations and empirical tests implied that the identified complex is H_{2}C_{2}⋯AgCCH rather than the original target of the experiment, H_{2}C_{2}⋯AgF. This deduction is now supported by evidence acquired through experiments exploiting ^{13}C-enriched isotopic samples. Transitions have been assigned for the H_{2}^{13}C_{2}⋯Ag^{13}C^{13}CH isotopologue. Data acquired from each isotopologue allows determination of the rotational constants (B_{0}, C_{0}) and centrifugal distortion constant, Δ_{J} . The data are consistent with a T-shaped complex in which the Ag atom of AgCCH binds to electrons within the π-orbitals of ethyne. Preliminary determinations of bond lengths will be presented. Experiments are in progress to measure the spectra of deuterated isotopologues.